OpenStudy (anonymous):

The table shows the cost of a ski rental package for a given number of people. People Cost ($) 4 160 5 200 6 240 7 280 The rate of change is constant in each table. Find the rate of change. Explain what the rate of change means for the situation.

OpenStudy (anonymous):

what is the answer

OpenStudy (texaschic101):

the rate of change is the slope. So we can find the rate of change using the slope formula. slope = (y2 - y1) / (x2 - x1) (4,160) x1 = 4 and y1 = 160 (5,200) x2 = 5 and y2 = 200 now we sub slope = (200 - 160) / (5 - 4) slope = 40 The rate of change is 40. Basically this means, for every person added, $40 is added to the rental package.

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

Answer the questions below. Make sure to show your work when applicable. 1. Solve the absolute value equation. Check your solutions. | 5x + 13| = –7

OpenStudy (anonymous):

2. Simplify the expression below. 6n2-5n2+7n2

OpenStudy (texaschic101):

| 5x + 13 | = -7 5x + 13 = -7 -(5x + 13) = -7 5x = -7 - 13 -5x - 13 = -7 5x = -20 -5x = -7 + 13 x = -4 -5x = 6 x = -6/5 x = -4 and x = -6/5 ========================= 6n^2 - 5n^2 + 7n^2 -- just combine like terms 8n^2 (because 6 - 5 = 1, and 1 + 7 = 8) the n^2 stays the same

OpenStudy (texaschic101):

3. 8b + 6 = 54 4. 8b + 6 = 54 8b = 54 - 6 8b = 48 b = 6 each bracelet costs $6 5. 6y - 8 <= 10 6y <= 10 - 8 6y <= 2 y <= 2/6 which reduces to 1/3, so y <= 1/3 6. not sure 7. 30% of 70 (of means multiply) -- change percent to decimal (.30)(70) = 21 8. -5 - 8 = - 13 9. commutative property 10. 4t + tu^2 - u^3 (t = 2 and u = 1) 4(2) + (2)(1^2) - 1^3 8 + 2 - 1 = 9 11. | r + 3| >= 7 r + 3 >= 7 r + 3 <= -7 r >= 7 - 3 r <= -7 - 3 r >= 4 r <= -10 r >= 4 and r <= -10 12. what percent of 60 is 12 percent formula : is/of = %/100 now we sub 12/60 = x/100 (cross multiply) 60x = 1200 x = 1200/60 x = 20% 13. sq rt 169 = 13 (because 13^2 = 169) 14. 100 + 25x >= 700 25x >= 700 - 100 25x >= 600 x >= 600/25 x >= 24 Sharon must sell at least 24 items for her to reach $700 or more per week. 15. - (-7y + 12) (-1)(-7y) + (-1)(12) 7y + (-12) 7y - 12 16. 1/a = 6/18 (cross multiply) (6)(a) = (18)(1) 6a = 18 a = 18/6 a = 3 17. 8x - 12 = 4x + 24 -- subtract 4x from both sides 8x - 4x - 12 = 24 -- add 12 to both sides 8x - 4x = 24 + 12 -- combine like terms 4x = 36 -- divide by 4 x = 9 18. -6b > 42 4b > -4 b < 42/-6 b > -4/4 b < - 7 b > -1 solution is : b < -7 and b > -1 19. 6 more then the product of 8 and n 8n + 6 20. 45 = 3b + 69 45 - 69 = 3b -24 = 3b -24/3 = b -8 = b

OpenStudy (anonymous):

Graph y = | x | + 5